OpenStudy (anonymous):

three elevenths __ 27% (1 point) A. > B.< C.=

OpenStudy (anonymous):

a. > 3/11 = .27272727272727272727... 27% = .27

OpenStudy (xishem):

Converting three elevenths to a percentage we get...\[\frac{3}{11}=0.27272727 \rightarrow 0.272727 * 100=27.2727percent\]\[27.2727>27\]